# Open Your Ears to Take a Look: A State-of-the-Art Report on the Integration of Sonification and Visualization

### v1

@@ Abstract
The research communities studying visualization and sonification for data display and analysis share exceptionally similar goals, essentially making data of any kind interpretable to humans. One community does so by using visual representations of data, the other community does so by employing auditory (non-speech) representations of data. While the two communities have a lot in common, they developed mostly in parallel over the course of the last few decades. With this STAR, we discuss a collection of work that bridges the borders of the two communities, hence a collection of work that aims to integrate the two techniques to one form of audiovisual display, which we argue to be “more than the sum of the two.” We introduce and motivate a classification system applicable to such audiovisual displays and categorize a corpus of 57 academic publications that appeared between 2011 and 2023 in categories such as reading level, dataset type, or evaluation system, to mention a few. The corpus also enables a meta-analysis of the field, including regularly occurring design patterns such as type of visualization and sonification techniques, or the use of visual and auditory channels, and the analysis of a co-author network of the field which shows individual teams without much interconnection. The body of work covered in this STAR also relates to three adjacent topics: audiovisual monitoring, accessibility, and audiovisual data art. These three topics are discussed individually in addition to the systematically conducted part of this research. The findings of this report may be used by researchers from both fields to understand the potentials and challenges of such integrated designs, while inspiring them for future collaboration with experts from the respective other field.
volume: 000issue: 0
@@ 1 Introduction
Over the course of the last few decades, two research communities have developed largely in parallel: one studying data visualization and one studying sonification. While the visualization community is primarily interested in “the use of computer-supported, interactive, visual representations of abstract data to amplify cognition” [CMS99], the sonification community studies “the use of non-speech audio to convey information” [$*{}^{*}$99]. To this day, the communities seem to be largely disjunct, despite their shared goals promising fruitful collaboration. Both the theoretical cross-pollination [$*{}^{*}$23, CLR22], and the practical integration and combination of sonification and visualization offer potential for interesting research outcomes. Therefore, with this state-of-the-art report (STAR), we want to shed light on audiovisual display idioms that systematically integrate data visualization and sonification. Informed by the definition of the visual idiom by Munzner [Mun15], we think of an audiovisual display idiom as “a distinct approach to creating and manipulating audiovisual representations of data.”
@@
In our daily lives, we perceive our surroundings in an inherently multimodal way. We see and we hear, we read books and we listen to music. We use our senses to understand and explore the world around us. Although we are multisensorial beings, the predominant data analysis idioms are uni-modal. Sight and hearing are inherently different, with different strengths and challenges, and are most probably suitable for different approaches in data representation. The visual perception can look upon a visual representation in a non-linear fashion, and the use of different types of graphs, charts, and other visual formats, can reveal patterns, correlations, and trends in data that are often not as noticeable in numerical form. Visual representations of data can also be experienced as more engaging and memorable to a user compared to tables with numbers [$*{}^{*}$21]. Sonification, on the other hand, exploits the excellent ability of the human auditory system to recognize temporal changes and patterns. It is, therefore, useful when displaying complex patterns, such as changes in time, and warnings for immediate action. In real-time environments, sonification allows a controller to perceive information without the constant monitoring of visual displays. Sonification also enables the communication of data and information for visually impaired individuals [WM10].
@@
The visual perception has some challenges that can be supported by sonification, and similarly the auditory system has other challenges that in turn can be supported by visualization. Therefore, we believe that a well-designed audiovisual representation can be more than the mere sum of a visual and an auditory representation. To put this into practice, we want to refer to a typical real-world situation where the combination of visual and auditory inputs aid us in reaching a more informed conclusion: imagine rain falling outside of a window. It often is hard to correctly estimate the density of the rain by just looking outside of a closed window. It is also not easy to estimate the amount of rain when only listening to it with your eyes closed. It is the holistic audiovisual perspective, that perceptually integrates both of our senses, that allows us to best determine whether we should use an umbrella or even stay indoors.
@@
Inspired by the capabilities of the human visual and auditory systems, and the possibility to integrate visualization and sonification, this STAR covers academic contributions from both the visualization and the sonification communities that blend sonification and visualization within the context of data exploration and data presentation. We hope this STAR will help both visualization and sonification researchers realize the potential of such combinations and foster future collaborations between two often disjunct communities.
@@ 1.1 Sonification Background
This STAR being published at a visualization venue calls for a brief introduction to the sonification techniques that are part of our data (for a more comprehensive description of the different techniques see the Sonification Handbook [HHN11]). Audification is a technique for representing sequences of data values that are transformed into sound. This sound is a translation of the data values into the audible domain, in terms of frequency and loudness. Audification typically represents a sequence of data as an audio waveform, that then can be processed with various signal processing techniques to further assess data features. Parameter mapping sonification involves the association of data values with auditory parameters such as pitch or loudness. The sound is dependent on the mapping function between the data and the auditory parameter and the nature of the mapping, for example being linear to linear or linear to exponential. Model-based sonification is a general term for various sonification techniques that make use of dynamic models of changes in a system over time. The sound is dependent on a sonification model based on the dynamics of a system that generates and changes the sound in relation to excitations caused by user interaction. Earcons are short distinctive sounds that are used to represent specific events or convey information. These sounds are often synthesized tones or sound patterns and can be described as designed or composed sound symbols. This presupposes an understanding and agreement about the meaning of the earcon before it can be beneficially used. Auditory icons are also short distinctive sounds but are sound snippets of real sound that are present in everyday life. This means that there is an association between the auditory icon and the event they are used to represent.
@@
A brief history of sonification: While the discipline of visualization has a relatively long history [Fri08], the research field of sonification is younger [Fry05], essentially starting with its first International Conference on Auditory Display in 1992. The proceedings of this first conference were published in 1994 in the book Auditory Display [Kra94]. When it was published, the book was also a reflection on the potentials of the newborn field of research and the newborn International Community For Auditory Display (ICAD). Early on, Barrass [Bar97] presented a task taxonomy for auditory displays, called TaDa!, which stands for Tasks and Data. The TaDa! taxonomy is especially relevant in the context of this STAR, as it is well-aligned with taxonomies from the visualization literature [Ber83, BM13, SNHS13, Shn96, YKSJ07], and it also functions as inspiration for the classification applied later in this STAR.
@@
The book Ecological Psychoacoustics, edited by Neuhoff in 2004 [Neu04], challenged many psychoacoustical studies (being the part of psychophysics that involves the scientific study of sound perception, traditionally conducted in controlled laboratory environments). Neuhoff promoted an ecological sound approach to sonification from a holistic perspective, which echoes the aims of the BELIV workshop series established in 2006 at the Advanced Visual Interfaces conference [BPS06]. Neuhoff’s intervention underscores the need to consider real-world contexts for transforming design principles and methodologies for auditory displays from theory to practice. This perspective emphasized integrating ecological factors in sonification to increase effectiveness and deepen the connection between auditory stimuli and real-world experiences.
@@
The sonification design space map was introduced in 2007 by deCampo [dC07], guiding a designer’s decision-making process of selecting an appropriate sonification technique for their task. The map coves a two-dimensional space between the number of data properties a designer intends to sonify, and the number of data points that is necessary for the different sonification techniques to be employed in an adequate manner. Retrospectively, another milestone within the sonification community was the introduction of the now widely accepted definition of sonification as a scientific technique for representing data, presented by Hermann [Her08] in 2008. Before the introduction of this definition, it was less clear where to draw the border between artistic and scientific mappings from data to sound (which brings to mind the discussion that data visualization is more than just pretty pictures). As it reflects our understanding of the term sonification, we want to refer to the full definition below:
@@
The Sonification Handbook [HHN11], published in 2011, provided the first general and overarching perspective on the field of sonification, discussing both theory and practice of sonification. The Handbook is still the most comprehensive collection of sonification work, and therefore its publication year in 2011 also marks the beginning of the time period considered in our STAR.
@@
Through the years, various design frameworks for sonification have been proposed. The design framework proposed by Barrass [Bar12], emphasizing the fusion of aesthetics and functionality to enhance the accessibility and meaningfulness of sonifications for a broader audience. The work done by Worrall in 2019 [Wor19b] formalized sonification techniques into a framework that also highlighted challenges and advantages of these sonification techniques as well as the importance of understanding the process and the choices that influence sound representation. The sonification design canvas, introduced by Lenzi in 2021 [Len21], is a contribution to constructing a more comprehensive design framework, aiming to integrate all aspects into a cohesive design tool. Despite these efforts, developing a comprehensive protocol that systematically considers end-users at each stage of the design process has yet to be achieved.
@@
More than 30 years after the beginning of systematic sonification research, we saw a considerable number of theoretical contributions to the field [Kra94, $*{}^{*}$99, VH06, dC07, NW08, HHN11, GH12, Sup12, Nee19, Neu19, Wor19b, Len21], nevertheless, explicit work integrating sonification and visualization theory is rare. In an attempt to find a common language and therefore build a theoretical bridge between the visualization and the sonification communities, Enge et al. [$*{}^{*}$23] introduced three theoretical constructs meant to formally describe audiovisual display idioms. They defined the “auditory mark” inspired by the visual mark, the “auditory channel” inspired by the visual channel, and the “substrate of sonification” inspired the spatial substrate that is employed in visualization theory [Ber83, CMS99]. The definition of time as the substrate of sonification allows a description of sonification designs with auditory marks being placed in time, with data encoded into their auditory appearance using auditory channels such as pitch of loudness. These definitions allow the a high level discussion and categorization of both the visual and the auditory part of an audiovisual display idiom. The theoretical constructs proved useful so that we adopted the term of auditory channels for our classification in this STAR as well.
1.2 Related Work and Surveys
@@
In general, systematic state-of-the-art reports are less established in the sonification community. A rare exception is the “systematic review of mapping strategies for the sonification of physical quantities” by Dubus and Bresin [DB13], which, however, just covers sonification-only contributions. Much earlier, in 2001, Walker and Lane [WL01] provided a website enabling researchers to search for sonification mappings that have been used in scientifically evaluated designs. Unfortunately, the website is no longer available. Another more timely and exhaustive exploration of sonification literature emerges in Andreopoulou and Goudarzi’s 2021 publication [AG21]. The authors reviewed 456 papers from the International Conference on Auditory Display proceedings. This incisive analysis exposes compelling trends, ranging from the sonification domains to the diverse publication venues. The report reveals linguistic trends and explores the balance between research and artistic contributions. In addition, it illuminates the landscape of tools, methodologies, and evaluation practices that have led to sonification’s multifaceted evolution. Marking the beginning a sociocultural reflection within the sonification community, in 2017, Andreopoulou and Goudarzi [AG17] also studied the “representation of female researchers and artists in the conferences of the International Community for Auditory Display (ICAD)”. Their findings showed that only about $18\%$ of ICAD papers were co-authored by women, with stagnant numbers between the years 1994 and 2016.
@@
With respect to combinations of visualization and sonification, Caiola et al. [CLR22] recently presented an analysis of visual and auditory channels commonly used in audiovisual display idioms. Their survey includes combined idioms that map data attributes redundantly to both a visual channel (such as position) and an auditory channel (such as pitch). Analyzed works stem from the Sonification Archive (described below) and a Google keyword search using sonification-related terms exclusively. (Sonification Archive), widely known in the sonification community, is a curated collection of sonification designs, often related to other modes of representation, such as visualization. The sonification archive holds both artistic and academic contributions, as well as designs from data journalism.

### v2

# Open Your Ears and Take a Look: A State-of-the-Art Report on the Integration of Sonification and Visualization

@@ Abstract
The research communities studying visualization and sonification for data display and analysis share exceptionally similar goals, essentially making data of any kind interpretable to humans. One community does so by using visual representations of data, and the other community employs auditory (non-speech) representations of data. While the two communities have a lot in common, they developed mostly in parallel over the course of the last few decades. With this STAR, we discuss a collection of work that bridges the borders of the two communities, hence a collection of work that aims to integrate the two techniques into one form of audiovisual display, which we argue to be “more than the sum of the two.” We introduce and motivate a classification system applicable to such audiovisual displays and categorize a corpus of 57 academic publications that appeared between 2011 and 2023 in categories such as reading level, dataset type, or evaluation system, to mention a few. The corpus also enables a meta-analysis of the field, including regularly occurring design patterns such as type of visualization and sonification techniques, or the use of visual and auditory channels, showing an overall diverse field with different designs. An analysis of a co-author network of the field shows individual teams without many interconnections. The body of work covered in this STAR also relates to three adjacent topics: audiovisual monitoring, accessibility, and audiovisual data art. These three topics are discussed individually in addition to the systematically conducted part of this research. The findings of this report may be used by researchers from both fields to understand the potentials and challenges of such integrated designs while hopefully inspiring them to collaborate with experts from the respective other field.
volume: 43issue: 3
@@ 1 Introduction
Over the course of the last few decades, two research communities have developed largely in parallel: one studying data visualization and one studying sonification. While the visualization community is primarily interested in “the use of computer-supported, interactive, visual representations of abstract data to amplify cognition” [CMS99], the sonification community studies “the use of non-speech audio to convey information” [KWB99]. To this day, the communities seem to be largely disjoint, despite their shared goals promising fruitful collaboration. Both the theoretical cross-pollination [ERI23, CLR22] and the practical integration and combination of sonification and visualization offer potential for interesting research outcomes. Therefore, with this state-of-the-art report (STAR), we want to shed light on audiovisual display idioms that systematically integrate data visualization and sonification. Informed by the definition of the visual idiom by Munzner [Mun15], we think of an audiovisual display idiom as “a distinct approach to creating and manipulating audiovisual representations of data.”
@@
In our daily lives, we perceive our surroundings in inherently multimodal and transmodal ways. We see, and we hear, we read books, and we listen to music. We use our senses to understand and explore the world around us. Although we are multisensorial beings, the predominant data analysis idioms are unimodal, often using visualization only. Sight and hearing are inherently different, with different strengths and challenges, and are most probably suitable for different approaches in data representation. Human visual perception can look upon a visual representation in a non-linear fashion, and the use of different types of graphs, charts, and other visual formats can reveal patterns, correlations, and trends in data that are often not as noticeable in numerical form. Visual representations of data can also be experienced as more engaging and memorable to a user compared to tables with numbers [FPS21]. Sonification, on the other hand, exploits the excellent ability of the human auditory system to recognize temporal changes and patterns. It is, therefore, useful when displaying complex patterns, such as changes in time and warnings for immediate action. In real-time environments, sonification allows a controller to perceive information without the constant monitoring of visual displays. Sonification also enables the communication of data and information for visually impaired individuals [WM10].
@@
The human visual perception has some challenges that can be supported by sonification, and similarly, the auditory system has other challenges that, in turn, can be supported by visualization. Therefore, we believe that a well-designed audiovisual representation can be more than the mere sum of a visual and an auditory representation. To put this into practice, we refer to a typical real-world situation where the combination of visual and auditory inputs aids us in reaching a more informed conclusion: imagine rain falling outside of a window. It is often difficult to correctly estimate the density of rain by just looking outside a closed window. It is also not easy to estimate the amount of rain when only listening to it with your eyes closed. It is the holistic audiovisual perspective, perceptually integrating both of our senses, that allows us to best determine whether we should use an umbrella or even stay indoors.
@@
Inspired by the capabilities of the human visual and auditory systems and the possibility of integrating visualization and sonification, this STAR covers academic contributions from both the visualization and the sonification communities that blend sonification and visualization within the context of data exploration and data presentation. We hope this STAR will help both visualization and sonification researchers realize the potential of such combinations and foster future collaborations between two often disjoint communities.
@@ 1.1 Sonification Background
This STAR being published at a visualization venue calls for a brief introduction to the sonification techniques that are part of our data (for a more comprehensive description of the different techniques, see the Sonification Handbook [HHN11]). Overall, we can distinguish the five main techniques of audification, parameter mapping sonification, model-based sonification, auditory icons, and earcons:
Audification is a technique to represent typically long sequences of data values (often time-series) by interpreting them as digital audio waveforms and directly playing them back over a loudspeaker. The resulting sound is a translation of the data values into the audible domain in terms of frequency and loudness. An example of audification is the playback of seismographic data with increased playback speed, such that the original low-frequency signal gets pitch-shifted into a range that is audible to humans. A design challenge for audification is the influence of the chosen playback speed on the salience of emerging auditory patterns that should inform a listener about their data.
Parameter mapping sonification (PMS) is a technique that involves the association of data values with auditory parameters such as pitch, timbre, and loudness. The technique is conceptually closest to many standard visualization techniques as it employs the direct mapping of data values to auditory channels of a carrying sound. If a visualization utilizes a visual mark to represent information by the mark’s position or its color, then a parameter mapping sonification utilizes a sound to carry information via its pitch, loudness, or other auditory channels. Similar to a visualization, such a sound is dependent on the mapping function between the data and the auditory channel as well as the nature of the mapping, for example, being linear to linear or linear to exponential. An open challenge in sonification research is the task-dependent and appropriate selection of auditory channels or the perceptual influence of one channel on another. In terms of visualization, this is comparable to the influence of the spatial size of a mark on our perception of its color.
The technique of model-based sonification (MBS) is inspired by the fact that most of the interactions with our physical environment result in an acoustic response. These acoustic responses, such as the sound of a drum being hit, inform us about the state of the object that we interact with (in the case of a drum, its tuning). Model-based sonification is a general term for sonification techniques that make use of dynamic models describing changes in a system over time. These dynamic models are “tuned” by the data that an analyst wants to explore. To listen to their data, a user is required to excite the model with an interaction, such as when a drummer needs to hit their instrument to hear its tuning. Another example is the excitation of a mass-spring model, where the mass and spring parameters are determined by high-dimensional data, defining the sound of the model when excited by a user. The technique is meant to foster exploratory data analysis, as the type and place of interaction strongly influence the acoustic response of the model. In this context, an ongoing research challenge is how to establish an intuitive relation between excitation modes (where and how hard to hit the drum) and typical interactions during exploratory data analysis, such as zooming or filtering of the data.
Also inspired by real-world sounds, auditory icons are short, distinctive sounds present in everyday life that can be compared to visual icons. This means that there is an inherent association between the auditory icon and the event they represent. A classic example of an auditory icon is the sound of a piece of paper being crumpled and thrown into a bin, representing deleting a file on a computer. A research challenge is the possibility of cultural differences between listeners, as they can lead to confusion or misinterpretation of auditory icons [JLSP15], similar to the way visual icons can be context-dependent.
Earcons, comparable to visual symbols, are short, distinctive sounds or melodies that are often used to represent specific events. These sounds are usually synthesized tones or sound patterns and can be described as designed or composed sound symbols. Since there is no inherent association with the real world, the meaning of the earcon needs to be learned before it can be beneficially used. Examples of such earcons are the sounds provided to pilots in the cockpit of an airplane, alerting them about events that require attention.
While other techniques for the sonification of data exist, these five are the most prominent ones. Within the scope of our STAR, we identified all five of these techniques, even though the vast majority of papers utilize parameter mapping sonification. Some of the papers also combine techniques (for example, parameter mapping sonification and auditory icons, see Table 1).
@@
A brief history of sonification: While the discipline of visualization has a relatively long history [Fri08], the research field of sonification is younger [Fry05]. In 1982, Bly submitted her PhD thesis Sound and computer information presentation [Bly82], where she suggested methods of encoding information into sound. Ten years later, the first International Conference on Auditory Display convened, which is regarded as the birth of the International Community For Auditory Display (ICAD). When its proceedings were published in 1994, the book Auditory Display [Kra94] was a reflection on the potential of the newborn field of research that encompasses sonification. Early on, Barrass [Bar97] presented a task taxonomy for auditory displays, called TaDa!, which stands for Tasks and Data. The TaDa! taxonomy is especially relevant in the context of this STAR, as it is well-aligned with taxonomies from the visualization literature [Ber83, BM13, SNHS13, Shn96, YKSJ07], and it also functions as inspiration for the classification applied later in this STAR.
@@
The book Ecological Psychoacoustics, edited by Neuhoff in 2004 [Neu04], challenged many psychoacoustical studies (which is the part of psychophysics that involves the scientific study of sound perception, traditionally conducted in controlled laboratory environments). Neuhoff promoted an ecological sound approach to sonification from a holistic perspective, which echoes the aims of the BELIV workshop series established in 2006 at the Advanced Visual Interfaces conference [BPS06]. Neuhoff’s intervention underscores the need to consider real-world contexts for transforming design principles and methodologies for auditory displays from theory to practice. This perspective emphasized integrating ecological factors in sonification to increase effectiveness and deepen the connection between auditory stimuli and real-world experiences.
@@
The sonification design space map was introduced in 2007 by deCampo [dC07], guiding a designer’s decision-making process of selecting an appropriate sonification technique for their task. The map creates a two-dimensional space between the number of data properties a designer intends to sonify and the number of data points that are necessary for the different sonification techniques to be employed adequately. Retrospectively, another milestone within the sonification community was the introduction of a now widely accepted definition of sonification as a scientific technique for representing data, presented by Hermann [Her08] in 2008. Before the introduction of this definition, it was less clear where to draw the border between artistic and scientific mappings from data to sound (which brings to mind the discussion that data visualization is more than just pretty pictures). As it reflects our understanding of the term sonification, we want to refer to the full definition below:
@@
The Sonification Handbook [HHN11], published in 2011, provided the first general and overarching perspective on the field of sonification, discussing both the theory and practice of sonification. The Handbook is still the most comprehensive collection of sonification work, and therefore, its publication year in 2011 also marks the beginning of the time period considered in our STAR.
@@
Over the years, various design frameworks for sonification have been proposed. The design framework proposed by Barrass [Bar12] emphasizes the fusion of aesthetics and functionality to improve the accessibility and meaningfulness of sonifications for a broader audience. The work of Worrall in 2019 [Wor19b] formalized sonification techniques into a framework that also highlighted the challenges and advantages of these sonification techniques, as well as the importance of understanding processes and choices that influence sound representation. The sonification design canvas, introduced by Lenzi in 2021 [Len21], is a contribution to the construction of a more comprehensive design framework, with the aim of integrating all aspects into a cohesive design tool. Despite these efforts, developing a comprehensive protocol that systematically considers end-users at each stage of the design process has yet to be achieved.
@@
More than 30 years after the beginning of systematic sonification research, we saw a considerable number of theoretical contributions to the field [Kra94, KWB99, VH06, dC07, NW08, HHN11, GH12, Sup12, Nee19, Neu19, Wor19b, Len21]. Explicit work integrating sonification and visualization theory is rare, but has been called for [RW10]. In an attempt to find a common language and, consequently, build a theoretical bridge between the visualization and the sonification communities, Enge et al. [ERI23] introduced three theoretical constructs to formally describe audiovisual display idioms. They defined the “auditory mark” inspired by the visual mark, the “auditory channel” inspired by the visual channel, and the “substrate of sonification” inspired by the spatial substrate that is used in visualization theory [Ber83, CMS99]. The definition of time as the substrate of sonification allows a description of sonification designs with auditory marks placed in time, with data encoded into their auditory appearance using auditory channels such as pitch or loudness. These definitions allow a high-level discussion and categorization of both the visual and the auditory part of an audiovisual display idiom. The theoretical constructs proved useful, so we adopted the term “auditory channels” for classification in this STAR as well.
1.2 Motivation and related work
Seminal visualization texts such as Wilkinson’s Grammar of Graphics [Wil05] and Spence’s textbook [Spe07] made clear statements that their understanding of visualization respectively graphics is not limited to vision but that data can be encoded for other sensory modalities such as sound. Already at the 6th IEEE Visualization Conference, Minghim and Forrest [MF95] postulated areas where sonification can help tackle visualization challenges such as adding complementary or redundant dimensions, natural mapping for time-oriented data, or improved memory of data. Some research agendas [TC05] suggest multimodal approaches so that one sensory modality can overcome problems that others may have. In 1990, Grinstein and colleagues [GS90, SBG90] presented an audiovisual interface for the exploration of multivariate data using icons and parameter mapping sonification. Also, several works presented at early visualization conferences have integrated sound into visualizations for surfaces, volumes, and fluid dynamics [MF95, LWS96, RN96, VG97, FBZ99]. The Data Mountain interface [RCL98] augmented its spatial document management space with auditory cues that indicated how many pages were moving. Published as early as 1989, Gaver presented the Sonic Finder, which was the auditory user interface used in Apple computers and coined the term “auditory icon.”
The auditory perception has an exceptional ability to detect temporal changes and patterns [GGB05, RG71]. Also, human hearing is capable of perceiving and distinguishing between several sounds simultaneously, at least between three auditory streams at the same time [SW13]. Another capability of our auditory perception is the possibility to detect and focus on events that spatially occur all around the listener, which can enable an information display to convey peripheral information to a user. Furthermore, the auditory system enables quick reactions when performing certain types of tasks due to the different processing times of the senses [JBKS15].
The combination of audio and visual data representations has shown to be advantageous, particularly for a number of applications. First, situations where the visual modality is busy with another task, such as monitoring [Ibe20, NB02], lend themselves beneficial for being complemented with sonification. Second, the combination of visual and auditory techniques has been shown to better facilitate learning [May14, SKS06], since it has the potential to increase working memory capacity and retention of information while also reducing cognitive load. Third, using sonification also shows benefits for data exploration. Flowers et al. [FBT97] published a seminal article in the sonification community that demonstrated that auditory scatter plots, where data is mapped to onset time and pitch of sounds instead of horizontal and vertical position of visual marks, can offer similar performance as visual representations.
Research has shown that visual and auditory perception is naturally integrated with each other, which can be observed, for example, with so-called crossmodal correspondences [Spe11, SDS23]. In our context, crossmodal correspondences describe a phenomenon where we perceive different visual and auditory stimuli as inherently related to each other. Such correspondences could provide an opportunity to use the strengths of the auditory modality to support and enhance visualizations, creating more effective and compelling representations of data [KBBG07]. Rosli and Cabrera [RC15], for example, have identified the potential of integrated designs to form a more concise, general representation of the data set compared to individual stimuli alone. Rubab et al. recently explored relationships between auditory channels and visual channels and suggested factors that impact their effectiveness [RYTW23].
Overall, the research described above suggests that the integration of the auditory modality has the potential to remedy challenges that exist for visual perception. Challenges, such as simultaneous brightness contrast [War19] or the Mach band phenomenon [LWP99], impact perception of visual representations [SGS18, ZTSS23]. It has been demonstrated that various auditory channels can be successfully linked and related to visual channels [FB18, WHT06, CH04], which could offer a way of substituting visual channels with auditory channels to address these challenges.
1.3 Related Surveys
@@
In general, systematic state-of-the-art reports are less established in the sonification community. A rare exception is the “systematic review of mapping strategies for the sonification of physical quantities” by Dubus and Bresin [DB13], which, however, just covers sonification-only contributions. Much earlier, in 2001, Walker and Lane [WL01] provided a website enabling researchers to search for sonification mappings that have been used in scientifically evaluated designs. Unfortunately, the website is no longer available. Another more timely and exhaustive exploration of sonification literature emerges in Andreopoulou and Goudarzi’s 2021 publication [AG21]. The authors reviewed 456 papers from the International Conference on Auditory Display proceedings. This incisive analysis exposes compelling trends, ranging from the sonification domains to the diverse publication venues. The report reveals linguistic trends and explores the balance between research and artistic contributions. In addition, it illuminates the landscape of tools, methodologies, and evaluation practices that have led to sonification’s multifaceted evolution. Marking the beginning of an important sociocultural reflection within the sonification community, in 2017, Andreopoulou and Goudarzi [AG17] also studied the “representation of female researchers and artists in the conferences of the International Community for Auditory Display (ICAD)”. Their findings showed that only about $18\%$ of ICAD papers were co-authored by women, with stagnant numbers between the years 1994 and 2016.
@@
With respect to combinations of visualization and sonification, Caiola et al. [CLR22] recently presented an analysis of visual and auditory channels commonly used in audiovisual display idioms. Their survey includes combined idioms that map data attributes redundantly to both a visual channel (such as position) and an auditory channel (such as pitch). Analyzed works stem from the Sonification Archive (described below) and a Google keyword search using sonification-related terms exclusively. The Sonification Archive, widely known in the sonification community, is a curated collection of sonification designs, often related to other modes of representation, such as visualization. The Sonification Archive holds both artistic and academic contributions, as well as designs from data journalism.